From the following data, calculate national income by income method
Items (Units of m)
Compensation of employees 800
Mixed income of self-employed 900
Net factor income from abroad - 50
Rent 350
Profit 600
Consumption of fixed capital 200
Net indirect taxes 250
Interest 450
Operating Surplus 1400
1
Expert's answer
2018-07-10T10:03:09-0400
According to income approach national income is:
NI =GNP - depreciation = GDP + net foreign income - depreciation = Rent + Interest + Profits + Wages + net foreign income - depreciation = 350 + 450 + 600 + 800 - 50 = 2150.